jeannie:

A 675-meter bike trail passes through forest and meadows. There are six legs of the trail that cut through meadows: two legs measure 24 3/4 meters, three legs measure 52.25 meters, and one leg measures 32 meters. The rest of the trail passes through the forest. How many meters of the trail pass through the forest? 238.25 meters 245 meters 436.75 meters 437.25 meters *edited by a moderator to include the complete question

jhonyy9:

try drawn it - may more easy understanding

Tranquility:

Hey there! @jeannie Welcome to QuestionCove! Your question seems incomplete "two legs measure meters, three legs measure 52.25 meters, and one leg measures 32 meters" What do the first two legs measure?

Tranquility:

I agree with Jhonny, drawing it out will help you understand the question better. Was there a diagram provided already? If not, we can make one. Are we supposed to find the value of those two legs? (Since they gave us the lengths of the other 4 legs)

jeannie:

Yes Ill try drawing it out. Thank you everyone

jhonyy9:

np. my pleasure

jhonyy9:

,, two legs measure ??? meters," from there signed with ??? you missed a number

Tranquility:

I think it's the missing number "x" that we need to solve for  perhaps something like this (note: not drawn to scale)

Gdeinward:

The only way to solve this is if the missing legs were equal in length. You should start by adding all the known measurements( the legs of the bike trail passing through meadow) and then subtracting it from the total known length(675). 675 = 156.75 + 32 + x +x 675 = 188.75 + x * 2 + Y 486.25 = x * 2 + Y However, this is where im stuck. With the given info it is impossible to solve without making guesses

Tranquility:

There is definitely information missing. Just read the question again. 675-meter bike trail passes through forest and meadows 6 legs of the trail cut through meadows remaining is in the forest so we need to at least know the value of all the legs. (this also makes my previous drawing incorrect)

Tranquility:

According to a previously asked (but unanswered) question: https://questioncove.com/updates/565ce006e4b0b09ed5dea614 A 675-meter bike trail passes through forest and meadows. There are six legs of the trail that cut through meadows: two legs measure 24 3/4 meters, three legs measure 52.25 meters, and one leg measures 32 meters. The rest of the trail passes through the forest. How many meters of the trail pass through the forest? 238.25 meters 245 meters 436.75 meters 437.25 meters

Tranquility:

So knowing that, in the equation that @Gdeinward made, we would plug in x = 24.75 and solve for 'y' which is how many meters of the trail passes through the forest

Gdeinward:

Thank you @Tranquility. So with the new info we can finish our equation: 486.25 = 24 3/4 * 2 + Y 24 3/4 = 24.75 because 3/4 = .75 486.25 = 24.75 * 2 + Y 486.25 = 49.5 + Y 486.25 - 49.5 = Y 436.75 = Y Therefore, the answer is C.